One leg of a right triangle is 2 ft more than twice the other leg. The hypotenuse is 1 ft. more than the longer leg. Find the length of the hypotenuse of the right triangle.
Answer by edjones(7569)

Let a=longer leg, b=shorter, c=hypotenuse.
a=2b+2
c=a+1
c=2b+2+1
a^2+b^2=c^2
(2b+2)^2+b^2=(2b+3)^2
4b^2+8b+4+b^2=4b^2+12b+9
b^2-4b-5=0 Subtract the right side from the left
(b-5)(b+1)=0
b=5'
a=12'
c=13'
.
Ed